OpenStudy (anonymous):

Julius has a garden that has dimensions of 12ft by 20ft. Julius needs more room to plant some peppers, and determines the garden will need a new area of 360sq ft. How much should he increase the length and width by. if he wants to increase them by the same amount?Round your answer to the nearest tenth. A.3.8Feet B.3.6Feet C.3.4Feet D.35.4Feet

OpenStudy (anonymous):

Please Help i will medal

OpenStudy (anonymous):

increase in area is equal to = new area - old area \[(12+x) \times (20+x) - (12 \times 20) = 360\]

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

none of the above?

OpenStudy (anonymous):

can you please tell me direct answer?

OpenStudy (anonymous):

I will medal you

OpenStudy (anonymous):

I dont need medal ..